# Single-Frequency Self-Alignment RF Resonant Beam for Information and Power Transfer

### v2

@@ I Introduction
To address these limitations and enhance high-efficiency transmission within the RF band, we propose the RF-based resonant beam system (RF-RBS). This system integrates RAAs and the resonance mechanism to facilitate self-alignment and achieve high-efficiency transmission for wireless power and information transfer. Figure 1 illustrates a smart home scenario where the RF-RBS is applied, where the base station (BS) is mounted on the ceiling, allowing IoT and mobile devices to be wirelessly charged and communicate simultaneously. The RF-RBS enhances power transfer efficiency and communication reliability while improving radiation safety and reducing EM interference, contributing to the advancement of IoT technologies.
@@
  • We propose a single-frequency RF-RBS architecture based on the resonance mechanism produced by a pair of RAAs to achieve high-efficiency and watt-level power transmission and support high-spectral-efficiency communication, which do not need additional beam steering devices or algorithms.
  • We developed an analytical model to represent the resonance mechanism of RF-RBS based on EM propagation theory and self-reproduction model, enabling the numerical determination of the system’s status and performance.
  • We analytically characterize the system convergence behavior by taking into account the noise introduced by the RAAs and other components across iterations. Specifically, RF-RBS can converge within a limited number of iterations in the presence of noise. Once the effective operating range is exceeded, the system’s radiated power decrease significantly to ensure EM radiation safety.
  • Compared to RD-BFS, RF-RBS achieves higher transmission efficiency due to its concentrated power distribution, resulting in up to a 16% improvement. Additionally, increasing the number of antennas in either the BS or the mobile target (MT) can extend the effective operating range of RF-RBS without extending the convergence time required for the system to position and track the MT.
@@ II-A System Design of RF-RBS
Figure 2 illustrates the transceiver design of the RF-RBS, including a BS and an MT, forming a mobile transmission channel.
@@ II-C Resonance Mechanism
To achieve resonance in RF-RBS, several conditions must be met: appropriate frequency, necessary boundary conditions, and a matched excitation source. The system’s operating frequency should be within a specific resonant frequency range, with all components tuned to a uniform frequency $f_{c}$. The boundary conditions must allow for proper reflection of EM waves, enabling multiple reflections to form a stable standing wave mode. In RF-RBS, RAAs enable EM wave transmission between BS and MT and ensure phase alignment through phase shifters. Finally, the excitation source must provide sufficient resonance power. In RF-RBS, the PAs supply the required energy, compensating for transmission, device, and output losses. Other devices, such as masers [35], which use stimulated radiation and resonance mechanisms for microwave amplification, are also viable options.
@@ III-A EM Waves Radiation and Propagation
To model multi-antenna signal transmission from $N$ transmitting antennas to $M$ receiving antennas, the transmission channel gain matrix $\mathbf{H}_{\text{c}}\in\mathbb{C}^{M\times N}$ is introduced, and for each element $h_{\text{c},n,m}$ in $\mathbf{H}_{\text{c}}$ is
@@
where $L_{nm}$ denotes the path length from the $n$-th transmitting antenna to the $m$-th receiving antenna. $G_{\text{t},n}$ and $G_{\text{r},m}$ are the gains of the transmitting and receiving antennas, respectively.
@@ III-F Communication Channel
where $\mathbf{w}\in\mathbb{C}^{1\times N}$ is the weight of the $n$-th antenna on the BS radiating to the selected antenna, and $F_{\text{d}}$ is the noise figure of the demodulation circuits.
@@ IV-D Power and Phase Distribution
Fig. 15 illustrates the phase distribution of the BS transmission and the MT reception when resonance is established. When the MT is located at $(0,0,2~{}\text{m})$ in Fig. 15(a), the phase distribution between the BS and the MT exhibits central symmetry, with isophase lines radiating outward from the center. This symmetrical distribution ensures that the focused signal is directed straight ahead, precisely aligning with the MT’s location. When the MT is located at $(0.5~{}\text{m},0,2~{}\text{m})$ in Fig. 15(b), the phase distribution becomes more complex. The isophase lines adjust, radiating outward from the new offset at both the BS and the MT. These lines also demonstrate a tendency to emanate from a specific coordinate point. By fitting the equivalent phase lines and calculating the center coordinates of the circles formed by the isophase lines, the results show that the distance between the antenna’s center and the MT’s offset along the $x$-axis are nearly equal.
@@ V Discussion
In the previous sections, we analyzed the point-to-point system. In this section, we discuss multi-target access within the RF-RBS. While an array composed of massive antennas can facilitate multi-directional transmission [54], the direction and distance of multiple MTs in RF-RBS can impact transmission loss, making power output to each MT uncontrollable [55]. To manage power outputs individually and enhance overall energy utilization, adopting an isolation scheme is essential. In RF-RBS, both Time Division Multiple Access (TDMA) and Frequency Division Multiple Access (FDMA) schemes are feasible.
@@ VI Conclusion
This paper demonstrated that RF-RBS operating in the RF band can achieve self-alignment and high-efficiency transmission based on the resonance mechanism. First, we introduced the RF-RBS architecture and the fundamentals of the RAA. Then, we established an analytical model to calculate power flow, resonance establishment, and communication performance. Numerical analysis revealed that the proposed RF-RBS consistently achieves self-alignment without beam control and provides higher transmission efficiency compared to RD-BFS, with improvements of up to 16%. Additionally, the results indicate the capability to transmit watt-level power and achieve 21 bps/Hz of downlink spectral efficiency in indoor settings. The more concentrated spatial power distribution of the RF-RBS enables high-efficiency transmission over long distances, highlighting the advantages of resonance in WPT and wireless communication applications. Future studies could investigate the system’s feasibility and performance under various scenarios and practical implementations. RF-RBS enables the IoT and mobile devices to operate without cables, presenting application prospects in smart homes, drones, and intelligent manufacturing. However, during implementation, several issues require further investigation, including i) developing multiple safeguards for high-power transmissions, ii) enhancing communication-perception integration, and iii) integrating Reconfigurable Intelligent Surfaces (RIS) to expand the working range.

### v5

@@ I Introduction
To address these limitations and enhance high-efficiency transmission within the RF band in the near field, we propose the RF-based resonant beam system (RF-RBS). This system integrates RAAs and the resonance mechanism to facilitate self-alignment and achieve high-efficiency transmission for wireless power and information transfer. Figure 1 illustrates a smart home scenario where the RF-RBS is applied, where the base station (BS) is mounted on the ceiling, allowing IoT and mobile devices to be wirelessly charged and communicate simultaneously. The RF-RBS enhances power transfer efficiency and communication reliability while improving radiation safety and reducing EM interference, contributing to the advancement of IoT technologies. Based on this conceptual framework, a simultaneous wireless information and power transfer (SWIPT) system is proposed with a focus on the design and performance analysis of a dual-frequency communication scheme [36]. However, the mechanism for achieving self-alignment and enhancing efficiency in RF-RBS remains unclear, and an analytical model for evaluating system performance under the three-dimensional (3D) movement of the mobile target (MT) is lacking. Additionally, a single-frequency design is preferred over the dual-frequency design for resonance establishment, as it reduces antenna configuration and hardware complexity. Thus, this work provides a detailed system design and develops models to elucidate the underlying mechanisms and analyze the performance of RF-RBS operating in a single frequency, particularly in terms of efficiency and mobility. Simulations incorporating the characteristics of practical components are conducted to validate the feasibility of the proposed RF-RBS approach.
@@
  • We propose a novel single-frequency RF-RBS architecture with low hardware complexity, leveraging the resonance mechanism and RAAs at both the BS and UD to achieve energy-concentrated and self-alignment transmission. The architecture requires the UD to participate in channel formation, enabling dynamic adaptation to arbitrary movements of the UD while maintaining high efficiency.
  • We establish a comprehensive theoretical model to analyze the proposed system based on EM propagation theory and a self-reproduction model, accounting for the noise introduced during the iterative process and establishing the corresponding transition matrices at each step. The model reveals the underlying mechanisms of resonance-driven transmission and demonstrates the system’s ability to overcome challenges such as sidelobe interference and beam tracking limitations.
  • Numerical evaluations show that the proposed system achieves high-efficiency transmission and high-spectral-efficiency communication in the near field without additional tracking control. Furthermore, increasing the number of antennas at either the BS or the MT extends the effective operating range of RF-RBS without increasing the convergence time required for positioning and tracking the MT.
@@ II-A System Design of RF-RBS
Figure 2 illustrates the transceiver design of the RF-RBS, including a BS and an MT operating in the near-field region, forming a mobile transmission channel.
@@ II-C Resonance Mechanism
To achieve resonance in RF-RBS, several conditions must be met: appropriate frequency, necessary boundary conditions, and a matched excitation source. The system’s operating frequency should be within a specific resonant frequency range, with all components tuned to a uniform single frequency $f_{c}$. The boundary conditions must allow for proper reflection of EM waves, enabling multiple reflections to form a stable standing wave mode. In RF-RBS, RAAs enable EM wave transmission between BS and MT and ensure phase alignment through phase shifters. Finally, the excitation source must provide sufficient resonance power. In RF-RBS, the PAs supply the required energy, compensating for transmission, device, and output losses. Other devices, such as masers [35], which use stimulated radiation and resonance mechanisms for microwave amplification, are also viable options.
The single-frequency design is specifically proposed in RF-RBS. Compared to the dual-frequency design [42], it eliminates the need for additional arrays with different element spacing and ensures the collinearity of the round-trip beams, further enhancing beam alignment accuracy and improving EM safety. However, a challenge arises from the phase delay difference between the BS and MT, which prevents the formation of standing waves in the mobile channel. As mentioned earlier, phase shifters in the BS are used to achieve the required phase alignment. Additionally, the design and manufacturing of high-gain, stable power amplifiers and phase shifters require further research.
@@ III-A EM Waves Radiation and Propagation
To model multi-antenna signal transmission from $N$ transmitting antennas to $M$ receiving antennas in the near-field region, the transmission channel gain matrix $\mathbf{H}_{\text{c}}\in\mathbb{C}^{M\times N}$ is introduced, and for each element $h_{\text{c},n,m}$ in $\mathbf{H}_{\text{c}}$ is
@@
where $L_{nm}$ denotes the path length from the $n$-th transmitting antenna to the $m$-th receiving antenna. $G_{\text{t},n}$ and $G_{\text{r},m}$ are the gains of the transmitting and receiving antennas, respectively. As the BS and MT operate in a near-field environment, EM wave propagation adheres to spherical wavefronts, requiring the phase differences across the antenna arrays to be rigorously calculated and accurately derived.
@@ III-F Communication Channel
where $F_{\text{d}}$ is the noise figure of the demodulation circuits.
@@ IV-D Power and Phase Distribution
Figure 15 illustrates the phase distribution of the BS transmission and the MT reception when resonance is established. When the MT is located at $(0,0,2~{}\text{m})$ in Fig. 15(a), the phase distribution between the BS and the MT exhibits central symmetry, with isophase lines radiating outward from the center. This symmetrical distribution ensures that the focused signal is directed straight ahead, precisely aligning with the MT’s location. When the MT is located at $(0.5~{}\text{m},0,2~{}\text{m})$ in Fig. 15(b), the phase distribution becomes more complex. The isophase lines adjust, radiating outward from the new offset at both the BS and the MT. These lines also demonstrate a tendency to emanate from a specific coordinate point. By fitting the equivalent phase lines and calculating the center coordinates of the circles formed by the isophase lines, the results show that the distance between the antenna’s center and the MT’s offset along the $x$-axis are nearly equal.
@@ V Discussion
In the previous sections, the point-to-point system is analyzed. In this section, we discuss multi-target access within the RF-RBS. While an array composed of massive antennas can facilitate multi-directional transmission [56], the direction and distance of multiple MTs in RF-RBS can impact transmission loss, making power output to each MT uncontrollable [57]. To manage power outputs individually and enhance overall energy utilization, adopting an isolation scheme is essential. In RF-RBS, both Time Division Multiple Access (TDMA) and Frequency Division Multiple Access (FDMA) schemes are feasible.
@@ VI Conclusion
This paper demonstrated that RF-RBS, operating in the RF band and within the near-field region, can achieve self-alignment and high-efficiency transmission based on the resonance mechanism. First, we introduced the RF-RBS architecture and the fundamentals of the RAA. Then, we established an analytical model to calculate power flow, resonance establishment, and communication performance. Numerical analysis revealed that the proposed RF-RBS consistently achieves self-alignment without beam control and provides higher transmission efficiency compared to RD-BFS, with improvements of up to 16%. Additionally, the results indicate the capability to transmit watt-level power and achieve 21 bps/Hz of downlink spectral efficiency in indoor settings. The more concentrated spatial power distribution of the RF-RBS enables high-efficiency transmission over long distances, highlighting the advantages of resonance in WPT and wireless communication applications. Future studies could investigate the system’s feasibility and performance under various scenarios and practical implementations. RF-RBS enables the IoT and mobile devices to operate without cables, presenting application prospects in smart homes, drones, and intelligent manufacturing. However, during implementation, several issues require further investigation, including i) developing multiple safeguards for high-power transmissions, ii) designing high-gain and stable PAs and phase shifters, iii) enhancing communication-perception integration, and iv) integrating Reconfigurable Intelligent Surfaces (RIS) to expand the working range.